40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2272     Entry time: Fri Jan 4 16:42:28 2019
Author: anchal 
Type: Summary 
Category: NoiseBudget 
Subject: Latest noisebudget update. 
  • Since we are working on temperature control of cavities, we worked out a hack to still take beat note spectrum meanwhile.

Method of measurement:

  • Spectrum is measured by SR785 at 800 points from 0 Hz to 6392 Hz.
  • No averaging is used and autorange is switched on.
  • Units of V_rms/sqrt(Hz) are used in the measurement.
  • I wrote a script, measBNnoise.py which is in Git, which continuously saves data from SR785 which is set as stated above.
  • Saving data takes from 90s to 120s, so data is being saved stroboscopically.
  • We are hoping that as the beatnote frequency is drifting slowly, Marconi will follow it almost always nicely except when it re-ranges.
  • So taking data in this way will mostly take data when Marconi is not re-ranging.
  • Data was taken for 12 hours in which total 381 measurements were made. 346 were filtered and used as described below.
  • Then I used another script I wrote, BNspectrum.py which does following:
    1) Reads all data files and computes the area under the curve. (We are using it as a figure of merit to filter out instances when Marconi was re-ranging)
    2) Throws away data sets which have this area above 2*sigma from the mean value of the whole set.
    3) Computes median at each frequency bin of remaining data set.
    4) Computes standard deviation from the median at each frequency bin.
    5) )Plots this data like iris (plot attached) and saves it in a .txt file to be used by noisebudget.ipynb later.

Present state of the experiment:

  • FSS on both paths has been upgraded to higher modulation frequencies (36 MHz on North, 37 MHz on South) (see PSL:2242 and PSL:2247)
  • PMC is installed on the South Path. (see PSL:2251)
  • See alignment details at PSL:2253.
  • ISS is not installed at this point.
  • Wideband NewFocus 1811 (DC-125MHz) RF Photodetector is used.
  • Temperature control on cavity shields was not on. They were put in a known almost stable heating state of 0.3871W Common and 0.7742W differential heating.

Conclusions:

  • This beatnote is as good as Feb 2018 measurement only.
  • In my opinion, ISS will help in reducing the noise further.
  • Presently, I have not optimized polarization of light going into EOM according to amplitude modulation produced by it. This in absence of ISS must be putting in more noise then we think.
  • Lens positions have not been optimized and are at there eyeballed positions where I put them after ala mode calculations. They might be off in order of cms.
  • The older elliptical filters in FSS boxes have not been replaced yet.
  • I'm not sure how big this is a factor, but in absence of temperature control, the frequency must be picking up more noise through drift in cavity lengths due to uncontrolled heating.
  • We are not using lower noise narrowband RF detector as we are presently not controlling BN frequency.
  • I also think we should have had taken data for a longer time. I will do this over the weekend.
  • I'll over the weekend write a script so that this hacked measurement will start happening every weekend. This way we can see any changes in our measurement from new things we work on during the week.
Attachment 1: Jan03_2019_BNnoise__Iris.pdf  5.540 MB  | Hide | Hide all
Jan03_2019_BNnoise__Iris.pdf
Attachment 2: 20190104_171447noiseBudget.pdf  929 kB  | Hide | Hide all
20190104_171447noiseBudget.pdf
ELOG V3.1.3-